# SpeechCap: Leveraging Playful Impact Captions to Facilitate Interpersonal Communication in Social Virtual Reality

### v1

@@ 3.1.2. Content Analysis
Three authors collaboratively conducted content analysis on the video collection to build the design space. First, three authors randomly selected a subset of videos (N=10) to individually code the visual elements and behaviors of the impact captions in the videos, with the aim of determining initial dimensions. Next, three authors sat together to reflect and discuss their initial insights and ultimately generated a consensus codebook. Finally, three authors reviewed and analyzed the complete data set. Consequently, we categorized the impact caption design space into visual space and interaction space, in which visual space focuses on textual and non-textual elements (Figure 2), and interaction space (Figure 3) relates to motions, physicalization, and embodied interactive operations.
3.2. Visual Design Space of Textual Elements
Textual elements are normally the primary visual component in an impact caption, including words, phrases, sentences, and symbols with linguistic meaning. In TV shows and videos, the content of textual elements comes from different sources such as narrative voiceover, conversations, onomatopoetic words, and commentary with ironic or entertaining implications added during the post-editing process of TV show production  (Sasamoto, 2014).
Typeface (font type), size, and color are three primary visual dimensions for designing textual elements. Depending on the contexts and purposes, all of these dimensions can be used to encode the importance of information (e.g., highlighting keywords) or represent a character’s emotion (e.g., explicitly showing moods).
Figure 2. Visual Design Space of Impact Captions. Impact captions consist of textual and non-textual visual elements. Textual elements include the typeface, color, and size of texts. Non-textual elements include emoji, ornament, and speech bubble.
@@ 3.2.1. Typeface
Typeface denotes the fundamental visual style of text and affects the reader’s perceptions and emotional responses  (Bianchi et al., 2021; Amare and Manning, 2012). Impact captions in videos widely utilize typeface to make impressions for audiences  (Sasamoto et al., 2021). However, there is no strict standard that defines which typeface should be used under which circumstance. In our design space, we summarized the empirical rules based on content analysis on TV show videos.
A key perspective of the typeface used in the impact caption is the regularity of word characters’ strokes. Artistic typefaces with irregular strokes are usually used to exaggerate characters’ moods or to emphasize surprising situations. The rich shapes of the strokes allow the typeface to match changing emotions and vivacious performance. Relatively, formal and regular typefaces (e.g., serif fonts) commonly appear in impact captions that are used as annotations with neutral information. This kind of caption provides additional knowledge to the original video content and should be optimized for viewers to read. In addition, the visual style of the texts can also be designed to fit the overall visual theme of the video in TV shows. In other words, various deeply customized typefaces were created for impact captions in current TV shows beyond standard typefaces used in common computer interfaces such as desktop and mobile phones.
@@ 3.2.2. Size
The size of captions is another dimension commonly used to encode semantic meanings. Size can be used for several purposes, depending on the creator’s design intent and the overall atmosphere and theme of the context. Texts with changeable sizes attract attention more effectively, making the subtitles stand out on the screen and piquing viewers’ interest. The use of larger fonts helps to emphasize key information, punchlines, or significant events. Additionally, larger fonts and flashy designs enhance the show’s entertainment value, contributing to a lighthearted and enjoyable experience for the viewers. Text size can also help create a specific atmosphere, with larger fonts conveying liveliness and excitement, while smaller fonts may evoke warmth and intimacy. In terms of readability, larger text is easier to read on television screens, ensuring that the information is easily accessible to the audience. Additionally, highly abstract semantic meanings, such as strength, effectiveness, and severity, can be encoded by text size in impact captions.
@@ 3.2.3. Color
In TV shows and videos, the color design of impact captions aims to create the show’s atmosphere, emphasize key information, and attract the audience’s attention. Thus, the color of impact captions usually features vibrant and eye-catching styles. For example, bright colors are usually used to add liveliness and appeal to entertainment-oriented shows, making captions more noticeable on the screen.
Besides, the color of impact captions is also used to imply the moods of characters, which reflects the color-emotion associations in human perceptions and self-expressions  (Plutchik and Kellerman, 2013; Hanada, 2018). Empirical evidence shows that bright and highly-saturated colors tend to associate with positive feelings, while dark and low-saturated colors may associate with negative feelings  (Wilms and Oberfeld, 2018; Hanada, 2018; Plutchik and Kellerman, 2013; Chen et al., 2021). However, there are no standard rules to tell which mood should be represented by which color. Because the color-emotion associations are not strict one-to-one mapping rules. A single color may reflect more than one type of emotion and vice versa  (Hanada, 2018; Wilms and Oberfeld, 2018). Moreover, the perceptions of emotion conveyed by the same color also vary among different groups of people  (Wilms and Oberfeld, 2018; Hanada, 2018; Chen et al., 2021).
Nevertheless, the color-emotion associations still highlight the potential of using color to convey emotional information for interpersonal communication bi-directionally. Color can help users receive emotional information from others and express emotional feelings of themselves.
3.3. Visual Design Space for Non-verbal Elements
Beside textual elements, several non-verbal elements, such as typographic ornaments, speech balloons, emojis (emoticons), are involved in impact caption. Those elements provide visual cues to convey invisible information, enhance expressiveness, and leverage engaging communication experiences.
@@ 3.3.1. Typographic ornaments
Typographic ornaments refer to decorative elements or symbols used to embellish and enhance the appearance of impact captions. They are typically small, non-alphabetic characters that add artistic flair and visual interest to the text. Common examples of ornaments include swashes, which are elegant flourishes on letters often seen in decorative scripts; fleurons, floral or leaf-like designs used as separators or accents; bullets, various shapes used for lists and highlighting points; dingbats, pictorial symbols representing a wide range of subjects; and borders and rule lines for separating text. In current impact captions, the content of ornaments is flexible and can be deeply customized to fit text contexts.
@@ 3.3.2. Speech bubbles
A speech bubble (a.k.a. speech balloon, ornamental frame) is the decorative visual element around text as a container of impact caption. This concept comes from comics and graphic novels and is used to represent the speech or thoughts of characters  (Aoki et al., 2022). It is a rounded or elliptical shape with a tail or pointer that connects to the character’s mouth, indicating the source of the speech. Inside the bubble, the character’s dialogue or thoughts are written in text to convey their words or inner monologue. In addition to speech, a speech bubble can also represent a character’s inner thoughts, visualize loud or intense sounds, and provide narrative captions (additional narration or context to the story). Speech balloons play a crucial role in visual storytelling, as they allow audiences to understand and follow the characters’ conversations and emotions. Different styles of speech balloons can be used to depict various emotions or tones of speech.
@@ 3.3.3. Emojis and Emoticons
Emojis are actual graphical images that depict a wide range of subjects, including facial expressions, animals, food, weather, etc. Similarly, an emoticon is a combination of characters, usually made from keyboard symbols, that represents a facial expression or an emotion  (Lo, 2008). Emoticons are commonly used in text-based communication, such as texting, messaging, social media posts, and online forums, to convey feelings or reactions that may not be easily expressed through text alone. Emojis and emoticons can be integrated into impact captions to enhance the emotional tone, clarity, and expressiveness.
3.4. Interactivity Design Space
When bringing impact captions from video-based media to VR, a significant change is that impact captions can be “alive” beyond unchangeable visual appearances and motions. In virtual reality, impact captions can be made interactive to respond to user input, including voice and embodied actions such as gestures and body movements. By playing with these virtual textual-visual elements, users can make fun, express themselves, and interact with others.
Figure 3. Interactivity Design Space of Impact Captions. The design space include three dimensions: Physicalization, Motion, and Interaction. With Physicalization, impact captions can appear like physical objects to be affected by the gravity with mass, have velocity for movement, and take spaces with volume. With Motion, impact captions appear to be “alive” and be responsive to user actions. Interaction describes how users can play with impact captions using embodied interaction.
@@ 3.4. Interactivity Design Space
3.4.1. Physicalization
Using the physical engine provided by current VR technology, impact captions can be endowed with various physical effects. These physical effects allow impact captions to behave like objects in the real world, helping users establish perception beyond language. For impact captions with specific meanings, physical sensations can be employed to enhance the semantics of the text. Our design space includes two of the most common physical quantities: mass and velocity Figure 3. Mass directly relates to people’s perception of “light” and “heavy”. When impact captions are assigned mass, gravity effects can be applied to them so that they would naturally convey the degree of abstraction in terms of weight. For example, assigning gravitational acceleration to an impact caption with the content “heavy-hearted” results in a dropping motion, physically conveying the feeling of “heaviness”. The concept of velocity is related to time. When adjusting the motion speed of impact captions, it can enhance the sense of urgency in expressing the textual content.
@@ 3.4.2. Motion
Motions make impact captions alive and attractive by endowing changing appearances. Motions can be independent (i.e., only affecting an impact caption itself) or dependent (i.e., relying on other objects to take effect). In our design space, we include three commonly observed independent animated motions: shivering, blinking, and explosion Figure 3. When shivering, primary visual components (i.e., textual and non-textual) of an impact caption will swing periodically. When blinking, the color of visual components (mainly text) could change rapidly and keep glowing. Explosion is an effect where an original caption splits into several pieces of fragments or replicas spread around the whole space. As for dependent motions, character (or avatar in VR) is the most common object with which impact captions can interact. Surrounding is a useful motion where impact captions move following an orbit. Surrounding impact captions usually represent the unique personalities or properties of avatars.
3.4.3. Interaction
Embodied interaction emphasizes the integration of the human body and physical actions, such as gestures  (Luo et al., 2024) and bodily movements  (Mueller et al., 2018), as integral components of interaction with digital technology  (Smith and Neff, 2018; Kirsh, 2013). Specifically, to support interpersonal communication using impact captions in VR, we recognized four common gestures (i.e., drag, drop, stretch, and throw) and one interaction with the virtual body (avatar) involved: attaching a caption with the avatar’s body. These embodied interactions can be used for either caption appearance customization or triggering particular actions. In this way, users can intuitively manipulate captions by their own hands rather than relying on a floating menu and clicking virtual buttons.
@@ 4. Co-designing an Impact-Caption-Inspired Communication Tool in Social VR
To concretize the design space to a feasible system design for our impact-caption-inspired approach, we conducted a co-design session with four HCI researchers and the four authors of this paper. From the session, we obtained primary directions and design goals for building the proof-of-concept system.
@@ 4.1. Procedure
The co-design procedure contains two sub-sessions. The first sub-session focused on brainstorming and open-ended discussion, and the second sub-session aimed at paper prototyping.
@@ 4.1.1. Participants
Besides the four authors of this paper, we recruited four extra HCI researchers from our university to our co-design session. During the process, the first author played the role of instructor to host the sessions and also participated in the activities to contribute ideas. The eight contributors are either developers or designers of digital products. They all have rich experiences in VR gaming and social VR (e.g., VRChat [^4] https://hello.vrchat.com/).
@@ 4.1.2. Sub-Session 1: Brainstorming and Discussion
In this sub-session, the main objective was to understand the scenarios, needs, and purposes of users’ interpersonal communication behaviors in social VR settings. This session was followed by initial brainstorming and design discussions. Participants were first asked to share their experiences and pain points to communicate with other VR players and discuss possible ways to address the issues. The instructor then presented several examples of impact captions (sourced from example videos collected during the design space investigation stage) to illustrate the concept and potential use cases of these impact captions. Next, the participants were asked to discuss their views on applying impact captions to social VR scenarios and brainstorm any desired ways to apply impact caption’s features to design communication tools in VR. The results of this session were eventually summarized by the instructor.
@@ 4.1.3. Sub-Session 2: Prototyping and Presentation
After brainstorming, the main objective was to continue the co-design process, concretizing ideas in the creation of low-fidelity prototypes. At the beginning, the instructor led a review of the results of the previous session and then informed the participants that they had 30 minutes to create their own prototype using white paper by drawing or handcrafting. Participants were encouraged to discuss their ideas with others and to think aloud during the session. However, participants still needed to create their own prototypes individually. When the time was up, each participant presented his/her prototype to the group, explaining the design concept, and discussing the prototypes.
We recorded the entire process of the co-design session with permissions from the contributors, transcribed it into text, and collected the prototypes they designed. Using the method proposed by Chen and Zhang  (Chen and Zhang, 2015), three of our authors conducted open coding and categorization of the design prototypes of the participants. In particular, components involved in the design of each prototype (such as input format, output results), information flow, context, and other visual elements (such as present style, text content, and colors) were coded by referring to existing work regarding Visual Captions  (Liu et al., 2023) and Augmented Presentation  (Liao et al., 2022). Finally, we revealed qualitative findings regarding how impact captions can support interpersonal communication and derived design goals for implementing a prototype system.
@@ 4.2. Findings
From the co-design activities, we identified three common patterns of using impact captions to facilitate interpersonal communication.
4.2.1. Enhancing Emotional Expressing
Emotions are abstract and inconspicuous. Impact captions could be an effective way to visualize and concretize human emotions as visible and touchable shapes in VR. This allows information senders in communication to share their emotional feelings in perceptible forms, so that the receivers could also easily understand the sender’s moods and intentions in the VR space.
The textual elements in the design space can encode various perspectives of emotional expressions. Text color is often used to represent positive or negative emotions in avatar dialogues. Generally, cool colors are used to convey more negative emotions, whereas warm colors represent more positive and enthusiastic emotions. For example, the word “indifference” could be rendered in deep blue, while “happy” is often depicted in bright orange or yellow colors. In addition to color, the typeface of text is also associated with emotions. More regular fonts are typically linked to serious and earnest emotions, whereas more artistic and irregular fonts convey a sense of openness and freedom. As for non-textual visual elements, emojis and emoticons are already commonly used elements for conveying people’s emotions in social media and instant messaging. Previous research has also found that different shapes of bubbles can carry different emotional information to express calm or intense expressions.
@@ 4.2.2. Highlighting Key Information in Speech
In conversations, impact captions can be used to highlight key information and, with their positioning or bubble identifiers, attribute them to the speakers. Typeface, color, and text size are useful dimensions for highlighting key information in a lency speech. Texts in distinctive colors or larger sizes are more likely to be read first by the recipient. Additionally, simpler and more regular fonts are more readable compared to overly complex artistic typefaces. In practical scenarios, multiple visual dimensions can be used individually or simultaneously to emphasize key information while aligning with the artistic style of the context.
@@ 4.2.3. Identifying Speakers
Although 3D audio and spatial sound effects allow users in social VR to trace the source of the sound and identify who is speaking, not everyone can accurately discern direction through simulated sound. Moreover, when multiple people speak simultaneously, the situation becomes more challenging. Impact captions, as visible virtual objects, are generated near the speaking avatars with a bubble tail pointing to the corresponding character in social VR. This allows participants to easily identify the current speaker through visual perception. Furthermore, when impact captions generated during conversations are allowed to persist in the environment for a prolonged period, newcomers can also see previous conversation content and gauge the level of activity of other participants based on the number of impact captions.

### v3

@@ 3.1.2. Content Analysis
Three authors collaboratively conducted content analysis on the video collection to build the design space. At the beginning, three authors started from a subset (N=10) of videos, 3 from the three popular TV shows and 7 randomly-selected online videos, to individually code the visual elements and behaviors of the impact captions in the videos, with the aim of determining initial dimensions. Through a round of discussion and merging on their initial insights, three authors ultimately determined the dimensions in the design space, including textual visual dimensions (i.e., typeface, size and color), non-textual visual dimensions (i.e., ornament, speech bubble and emoji), and interactions (i.e., physicalization, motion, embodied interaction).
Using the determined dimensions as a codebook, each of the three authors took 13 videos (i.e., one third of the rest videos) to review. During the process, the reviewers also had multiple rounds of discussions on the possible values of some dimensions (e.g., spiky bubbles in speech bubble). As a result, we found that all of the 49 (i.e., 3 from the TV shows plus 46 online) videos manipulated both textual and non-textual visual dimensions to design impact captions. For the dimensions under the interaction category, 29 videos significant used motions, 13 videos had physicalization, and 15 made impact captions to interplay with human characters.
3.2. The Visual Design of Textual Elements
As the primary visual component, the textual elements of impact captions follow the principles of general typographic design  (Carter et al., 2011), in which typeface, size, and color are three main visual dimensions.
Figure 2. The Visual Design Space of Impact Captions. Two impact captions, a “No Way” and a “Shocked Face” (translated from Chinese), demonstrate the textual and non-textual elements with relevant dimensions that form the visual design space. For textual elements, the typeface, the color, and the size of texts are three major dimensions. Non-textual elements include emoji, ornament, and speech bubble.
@@ 3.2.1. Typeface
Typeface denotes the fundamental visual style of each letters in a piece of text  (Carter et al., 2011). It can affect the perceptions and emotional responses of the readers  (Bianchi et al., 2021; Amare and Manning, 2012). Impact captions widely utilize typeface to make impressions for the audiences  (Sasamoto et al., 2021). Although there is no strict standard that defines which typeface should be used under which circumstance, we found that the regularity of strokes can relate to the content of impact captions. Specifically, artistic typefaces with irregular strokes are usually used to highlight characters’ moods or to emphasize surprising situations, while formal and regular typefaces (e.g., serif fonts) are used for presenting neutral and factual information.
@@ 3.2.2. Size
The size of the textual elements is a powerful dimension used for several purposes. First, impact captions with changeable sizes attract attention, making the particular text content stand out on the screen to engage viewers. Second, larger texts helps enhance the readability and accessibility of video content. Third, size can convey liveliness atmosphere and the excitement of characters. Additionally, size can also encode some highly abstract semantic meanings, such as strength, effectiveness, and severity.
@@ 3.2.3. Color
Similar to size, color can be vibrant and eye-catching to attract the audience’s attention and highlight important information. In addition, color of the textual elements of impact captions is commonly used to imply the moods of characters in TV shows, which reflects the color-emotion associations in human perceptions  (Plutchik and Kellerman, 2013; Hanada, 2018). Although the perceptions of the emotion conveyed by the same color can vary among different people  (Wilms and Oberfeld, 2018; Hanada, 2018; Chen et al., 2021), the color-emotion associations still reveal the potential of color as a dimension to convey emotional information for bi-directional interpersonal communication.
3.3. The Visual Design of Non-textual Elements
We identified typographic ornaments, speech bubbles, and emojis & emoticons as three types of commonly used non-textual elements of impact captions through the content analysis.
@@ 3.3.1. Typographic ornaments
Typographic ornaments refer to the decorative elements or symbols used to embellish and enhance the appearance of impact captions. They are typically small, non-alphabetic characters that add artistic flair and visual interest to the textual elements. For example, the “erupting gas” ornaments strengthen the tone of saying “No Way” (Figure 2).
@@ 3.3.2. Speech bubbles
Speech bubbles (a.k.a. speech balloons, ornamental frames) refer to the background containers to display text inside. In comics, speech bubbles is widely used for presenting speech or thoughts and indicating the source of the speech or thoughts with a tail pointing to characters  (Cohn, 2013). As a non-textual visual element in impact captions, speech bubbles not only inherit the functionalities as they have in comics, but also convey non-verbal information such as the emotion of characters and the tone of speech through shape and color. This feature has been in other contexts like text messaging  (Aoki et al., 2022).
@@ 3.3.3. Emojis and Emoticons
Emojis and Emoticons, as non-textual glyphs that represent facial expressions  (Lo, 2008), are a special group of decorative elements used in impact captions to convey feelings or simulate human reactions that can not be easily expressed through texts alone. For example, the “amazed face” emoji in the lower impact caption in Figure 2 visually repeats the meaning of the text to enhance the impressiveness.
3.4. The Interaction Design of Impact Captions
Shifting from video to VR, impact captions are no longer post-edited visual effects but become “alive”—real-time rendering allows impact captions to lively interact with users, while the human-caption interactions in VR remain under-explored. Through the content analysis, we found that impact captions in TV shows are often designed to simulate physical properties and possess motion effects, or made to coordinate with the embodied actions of human characters. Inspired by these post-edited behaviors, we explored the interaction design space of impact captions for VR, considering Physicalization, Motion, and Embodied Interaction.
3.4.1. Physicalization
Physicalization allows virtual objects to simulate physical properties to create tangible sensations for users to enrich their interactions with digital user interfaces  (Sauvé et al., 2024; Hornecker and Buur, 2006). We found that impact captions can also appear like they have physical properties in TV show videos to convey non-verbal meanings. Three commonly used physical properties are mass, velocity, and volume (Figure 3 A). Mass implies the perception of weight and can be used to represent the feeling of “light” and “heavy” of abstract concepts such as the seriousness of speech. Velocity associates with time, indicating the sense of urgency. Volume relate to the size of text in impact captions but the overall volume of an impact caption can still convey the sense of “small” and “big” without words.
Figure 3. The Interaction Design of Impact Captions. The design space includes three dimensions: Physicalization, Motion, and Interaction. With Physicalization, impact captions can appear like physical objects to be affected by the gravity with mass, have velocity for movement, and take spaces with volume. With Motion, impact captions appear to be “alive” and be responsive to user actions. Interaction describes how users can play with impact captions using embodied interaction.
@@ 3.4.2. Motion
Motions make impact captions alive and attractive by endowing changing appearances. Four kinds of motions were found and included in our design space: shivering, blinking, and explosion (Figure 3 B). Shivering means that impact captions will swing periodically . Blinking makes the color of visual components (e.g., text or speech bubble) change rapidly and glowing. Explosion is an effect where an original caption splits into several pieces of fragments or replicas spreading around. Circling makes impact captions orbit an object or human characters.
3.4.3. Embodied Interaction
Embodied interaction emphasizes the integration of the human body and physical actions, such as gestures  (Luo et al., 2024) and bodily movements  (Mueller et al., 2018), as integral components of interaction with digital technology  (Smith and Neff, 2018; Kirsh, 2013). Inspired by the patterns of how impact captions interplay with characters in TV shows, we include four gestures (i.e., drag, drop, stretch, and throw) and a bodily movement (i.e., attach to avatar) in our design space. With these embodied interactions as triggers, users can intuitively manipulate and play with impact captions in VR.
@@ 4. Co-designing an Impact-Caption-Inspired Communication Tool in Social VR
To operationalize our impact-caption-inspired approach for interpersonal communication in social VR, we organized a co-design activity to obtain and refine the design goals for build a proof-of-concept system.
@@ 4.1. Procedure
The procedure consists of two sub-sessions. The first focused on brainstorming and open-ended discussion, and the second aimed at paper prototyping for the proof-of-concept system.
@@ 4.1.1. Participants
Besides the four authors of this paper, we recruited four extra HCI researchers from our university to this session. During the procedure, the first author played the role of instructor to host the sessions and also participated in the activities to contribute ideas. The eight contributors are either developers or designers of digital products. They all have rich experiences in VR gaming and social VR (e.g., VRChat [^4] https://hello.vrchat.com/).
@@ 4.1.2. Sub-session 1: Brainstorming and Discussion
In the first sub-session, the objective was to understand the scenarios, needs, and purposes of users’ interpersonal communication behaviors in social VR. This session was followed by an initial brainstorming and discussion activity in which the participants shared their own experiences of interpersonal communication in VR to identify the potential pain points and possible improvements. The instructor then presented several examples of impact captions (sourced from the sample videos for investigating the design space) to illustrate the concept and potential usages. With the inspiration of impact captions, next the participants presented their views on how the captions can be applied to improve the communication experience in social VR. The results of this session were recorded and summarized by the instructor.
@@ 4.1.3. Sub-session 2: Prototyping and Presentation
After the brainstorming and discussion, the goal of the second sub-session is to create low-fidelity prototypes so that the previous ideas can be visualized and demonstrated in more concrete forms. At the beginning of this sub-session, the instructor led a review of the results of the previous session and then informed the participants that they had 30 minutes to create their own prototype using white paper by drawing or handcrafting. Participants were encouraged to discuss their ideas with others and to think aloud during the session, while they still needed to create prototypes individually. After the creation was completed, the participants presented their prototypes to the group and developed open-ended discussions on the prototypes.
4.1.4. Data Analysis
We recorded the entire procedure of the co-design activity with the permissions of the participants, transcribed oral discussions into texts, and collected the prototypes they designed. Using the method proposed by Chen and Zhang  (Chen and Zhang, 2015), three authors worked as coders to conduct open coding and categorization of the design prototypes. In particular, components involved in the design of each prototype (such as expected input and output), information flow, context, and other visual elements (such as present style, text content, and colors) were coded by referring to related work about augmenting presentation and communications  (Liu et al., 2023; Liao et al., 2022; de Lacerda Pataca et al., 2024). Finally, we revealed qualitative findings regarding how impact captions can support interpersonal communication and derived design goals for a prototype system.
@@ 4.2. Findings
Through the co-design activity, we identified three aspects of interpersonal communication in which impact captions can help.
4.2.1. Enhancing Emotional Expression
Emotions are important in communication but sometimes it could be implicit and inconspicuous. Impact captions can be a way to materialize human emotions into visible and tangible shapes in social VR. With impact captions, information senders can share their emotional feelings in visible and touchable forms so that the receivers can perceive the sender’s moods and intentions with multi-modal perceptions beyond verbal signals and body movements.
@@ 4.2.2. Highlighting Key Information in Speech
Impact captions can be used to highlight key information in a conversation. Unlike normal captioning system that aims to display all the words in a speech, impact captions should be made for words that conveys key information. Additionally, typeface, color, and text size are useful dimensions to enhance the visual attractiveness of impact captions to make the key information visible and outstanding in a lengthy speech in VR.
@@ 4.2.3. Identifying Speakers
Although spatial audio allows users of social VR to trace the source of sounds and identify speakers, it still lacks accuracy and explicit signals, especially when multiple people speak simultaneously  (Yan et al., 2023). Impact captions, with their positioning or speech bubble identifiers, can help users to easily identify the source of speech or sound through visual cues. This can be a complement to the audio cues to enrich the communication experience in social VR.